40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Mon Apr 15 21:11:49 2019, Philip, Optics, Characterization, OMC(004): PZT testing for spare OMC 
    Reply  Tue Apr 16 11:36:36 2019, Koji, Optics, Characterization, OMC(004): PZT testing for spare OMC 9x
    Reply  Tue Apr 16 21:16:11 2019, Koji, Optics, Characterization, OMC(004): PZT testing for spare OMC 
Message ID: 336     Entry time: Mon Apr 15 21:11:49 2019     Reply to this: 337   342
Author: Philip 
Type: Optics 
Category: Characterization 
Subject: OMC(004): PZT testing for spare OMC 

[Koji, Philip]

Today we tested the functionality of the four remaining PZTs (11,12,13 and 22) .  Each PZT was placed within a collimated 500um beam.
Roughly half of the beam was blocked by the PZT. The PZT and a PD then acted as shadow sensor. Each PZT was tested with 0 and
150 V. The resulting power change then could be converted into a displacement of the PZT using the beam diameter.

The open light value for each of these tests was 3.25 V.

PZT 11:
0 V supply voltage     --> 1.717 V on PD
150 V supply voltage --> 1.709 V on PD
delta = 0.008 V

PZT 12:
0 V supply voltage     --> 1.716 V on PD
150 V supply voltage --> 1.709 V on PD
delta = 0.007 V

PZT 13:
0 V supply voltage     --> 1.702 V on PD
150 V supply voltage --> 1.694 V on PD
delta = 0.008 V

PZT 22:
0 V supply voltage     --> 1.770 V on PD
150 V supply voltage --> 1.762 V on PD
delta = 0.008 V

0.008 V --> 0.24% change in power on PD --> about  3.8 um displacement assuming no light which is blocked
by the PZT is hitting the PD.

 

We further started to drive all four PZTs over night with 100 V (half of their range) at 100 Hz.
We additionally display the impedance to ensure none of them degrades.

All four PZTs seem to be connected to Teflon coated wires. It needs to be checked if these
fulfill the vacuum compatibility requirements.

ELOG V3.1.3-